Consider the 2-dimensional phase space R^2. Find the most general canonical transformation of this phase space,of the form Q=Q(q,p)=f(q)+g(p)

P=P(q,p)=c[f(q)+h(p)] where f,g and h are differentiable real-valued functions


1
Expert's answer
2021-05-06T19:12:58-0400

Given equations are,

"Q=Q(q,p)=f(q)+g(p)"


"P=P(q,p)=c[f(q)+h(p)]"


"\dot{Q}=Q(p)\frac{dQ(q)}{dq}+Q(q)\frac{dQ(p)}{dq}"


"=f'(q)+g'(p)"


"\dot{P}=P(p)\frac{d(P(p))}{dp}+P(q)\frac{dP(q)}{dp}"


"=c[f'(q)+h'(p)]"